# Variational Scarring in Open Two-Dimensional Quantum Dots

### v1

@@ Abstract
Quantum chaos – sometimes referred to as quantum chaology [1, 2] – explores the transition from quantum states to classical behavior, a topic of intense debate since the advent of quantum mechanics [3, 4]. While classical mechanics exhibits true chaos, quantum systems display signatures of chaotic behavior in distinct ways [5, 6, 7, 8], such as characteristic spectral statistics, wavefunction scarring, and other non-ergodic phenomena. At the heart of this exploration lies the correspondence principle, which intimately linked to longstanding questions regarding measurement and decoherence in quantum mechanics [9, 10]. One proposed framework for understanding the emergence of classicality is quantum Darwinism, which posits that decoherence – by suppressing most quantum superpositions – yields environment-induced superselection (einselection). This process favors the survival of robust, pointer states that proliferate their information into the environment, while fragile quantum states are suppressed [11]. Intriguingly, these pointer states often exhibit quantum scarring, thus defying the expected ergodic behavior of quantum-chaotic systems [12].
@@
In general, open quantum dots (QDs) provide a versatile platform for investigating quantum scars, offering valuable insights into classical-quantum correspondence partially due to their mixed classical phase space structure. While a single scarred state is typically thought to be fragile against strong perturbations, such as system opening, the proliferation of scars can however results in significant robustness for the entire family of scarred states within a moderate energy window. Consequently, these scars can have a substantial effect on the transmission properties of quantum devices [29, 33, 43, 44]: notable transmission variations over small energy scales have been observed, with the fractal conductance fluctuations in quantum devices often attributed to their chaotic nature.
@@
In this Letter, we investigate the quantum transport of a two-dimensional (2D) device illustrated in Fig.1 in which the coupling to the reservoirs is mediated by a pair of quantum point contacts. We assume that the QD contains impurities introducing randomly positioned bumps within the confining potential, whereas a more controlled, localized disturbance is generated via a nanotip. In particular, the perturbation gives rise a recently discovered form of scarring within the closed QD [45, 46, 47, 48, 49, 50, 51]: these scars emerge in a synthesis of the quantum near-degeneracy in the unperturbed system and the localized nature of the perturbation. This novel birth mechanism [48] sets the variational scars apart from the conventional, Heller-type scarring [13, 17, 18], extending the concept of scarring into systems lacking suitable classical POs or when those orbits are too unstable to sustain conventional scarring. In other words, this novel origin allows scars to emerge and thrive under conditions that would typically inhibit conventional scarring. Furthermore, it highlights the fact that the near-degeneracies of a system can have a profound impact, not only the exact degeneracies associated to the continuous symmetries, as well as underscores the subtle role that disorder can play in a quantum device.
@@
In addition to the confining potential $V_{\textrm{ext}}$, recognizing that real quantum devices are often affected by disorder arising from impurities and imperfections – such as atoms or ions migrating into a QD system – we model the disorder potential as a sum of randomly distributed Gaussian bumps $V_{\text{imp}}=A_{n}\sum_{i}\exp(-\frac{|\bm{r}-\bm{r}_{i}|^{2}}{2\sigma_{n}^{2}})$, where $A_{n}$ and $\sigma_{n}$ represent the amplitude and width of the impurity bumps, respectively. This kind of disorder mode has been validated for a QD by density-functional theory [68, 69] and the diffusive quantum Monte-Carlo approach [62]. Moreover, we introduce a controllable localized perturbation stemming from a conducting nanotip for which a well-established approximation [70, 71, 72] is also the Gaussian profile $V_{\text{tip}}=A_{T}\exp({\frac{|\mathbf{r}-\mathbf{r}_{0}|^{2}}{2\sigma_{T}^{2}}})$, centered at the location $\bm{r}_{0}$, where $A_{T}$ and $\sigma_{T}$ denote its amplitude and width.
@@
As a reference, we first consider the case where there is no nanotip (i.e., $V_{T}=0$), and the only perturbation comes from the impurity bumps $V_{\textrm{imp}}$, with amplitude $A_{n}=2$ and width $\sigma_{n}\simeq 0.05$. In this scenario, we find that several states accommodate a BB scar. Due to the spatially localized nature of the perturbation, these scarred states are favored according to the variational principle [45, 48]: the BB scars shown in Fig.2(a) orient themselves to maximize or minimize the perturbation by coinciding with (or avoiding) as many bumps as possible. While the scars maintain a similar alignment across a broad energy range, controlling the scar direction is challenging due to the inherent randomness of the impurity perturbation. However, by introducing an additional, stronger bump (i.e., $V_{T}\neq 0$) via the nanotip ($A_{T}=12$ and $\sigma_{T}=0.1$), the BB scars align with the position of the nanotip, as seen in Fig.2(b), which provides a feasible strategy to design the orientation of the scars experimentally. Furthermore, it is also worth noting that the single perturbation caused by the tip can be sufficient to generate distinct scars [46].
@@
Figure 2: (a) Examples of eigenstates in a closed, perturbed quantum dot containing random impurities (characterized by amplitude $A_{n}=2$ and width $\sigma_{n}\simeq 0.05$). The positions of impurities are represented by green dots. (b) Examples of eigenstates with the same system and quantum numbers $n$ as in (a), but now including a nanotip generating a additional, more prominent potential bump (amplitude $A_{T}=12$, width $\sigma_{T}=0.1$), which is indicated as a red dot.
@@
Previous work [45] has demonstrated variational scars to impact the transport characteristics of a QD by enabling wavepackets to propagate through the perturbed system with higher fidelity than in the corresponding unperturbed system. However, the influence of these scars on the electrical transmission of an open quantum device has remained elusive. To address this question, we couple the QD to two leads, effectively opening the system. The leads are modeled as semi-infinite waveguides, with harmonic confinement in the transverse direction and free propagation along the longitudinal axis. This setup ensures smooth continuity and consistent boundary conditions at the interface with the scattering region. The coupling between the leads and the QD is characterized the rate operator $\Gamma$ that subsequently defines the coupling parameter $\mathcal{C}=\textrm{Tr}[\hat{\Gamma}(\omega)]/\Delta E$ where $\Delta E$ represents the numerically imposed energy spacing in the effectively continuous spectrum of the leads.. The transmission through the QD is then determined within the Landauer-Büttiker formalism [74, 75, 76] as
$\mathcal{T}(\omega)=\text{Tr}\Big{[}\hat{G}^{R}(\omega)\,\hat{\Gamma}(\omega)\,\hat{G}^{A}(\omega)\,\hat{\Gamma}(\omega)\Big{]},$ (2)
@@
$\rho(\bm{r},\omega)=-\text{Im}\big{[}\hat{G}^{R}(\bm{r},\omega)\big{]},$ (3)
@@
Fig.3 shows the transmission $\mathcal{T}(\omega)$ of the device as a function of energy $\omega$ for weak coupling $\mathcal{T}(\omega)$ between the leads and the QD. Like in the scarring analysis above, we compare the transmission with the nanotip to the situation with only impurity disorder $V_{\textrm{imp}}$. Notably, we observe significantly higher transmission peaks in the nanotip scenario [peak labeled (a) in Fig.3] compared to the noise-only instance [peak labeled (b) in Fig.3]. We attribute this boost in transmission to the presence of BB scarring, where the noise-only case lacks the coherent pathways formed by the lead-oriented scars. This interpretation is confirmed by the LDOS, which reveals a pattern at the transmission peak energy [Fig.3(a)] corresponding to a BB scar of the closed QD, a feature absent in the nanotip-free benchmark [Fig.3(b)].
@@
Figure 3: Logarithm of transmission vs. probe energy within bias voltage 23–24 for noise-only (blue) and nanotip+noise (orange) in weak coupling limit $\mathcal{C}\propto 1$. Peaks are labeled as (a) for noise-only and (b) for nanotip+noise. Logarithmic scale is used due to the wide range of transmission values. Below, normalized LDOS at corresponding peak energies are shown for (a) noise-only and (b) nanotip+noise, illustrating distinct bouncing-ball scar pattern induced by the tip. Panels (c)-(e) depict normalized LDOS of nanotip+noise at the same peak energy (23.350) in weakly coupled leads, now under stronger coupling strengths $\mathcal{C}$ : (c) $\propto 10^{2}$, (d) $\propto 10^{4}$, and (e) $\propto 10^{5}$.
@@
On the other hand, panels (c)-(e) of Fig.3 present the LDOS for the nanotip case at the same energy of 23.350, corresponding to the strongest transmission peak in the weak coupling limit. These cases correspond to three different coupling strengths, $\mathcal{C}\propto 10^{2},10^{4},10^{5}$, respectively. As the coupling strength increases, the LDOS pattern that initially contains the scar gradually becomes more smeared and chaotic. This reflects the fact that stronger coupling weakens the influence of the individual scar on electron transport by broadening the QD energy levels. To explore the strong coupling limit ($\mathcal{C}\gg 1$) further, we focus on the same energy range where the nanotip-induced transport enhancements occur under weak coupling. Therefore, Fig.4 shows the transmission for two cases of strong coupling ($\mathcal{C}\propto 10^{2}$ and $\mathcal{C}\propto 10^{5}$). As illustrated in Fig.4(a) and Fig.4(b), the level broadening caused by the leads disrupts the scar patterns seen in the weak coupling regime, replacing them with more extended density distributions. Furthermore, this pronounced level broadening alters the line shape of the transmission peaks. In the weak coupling regime, the QD exhibits asymmetrical, Fano-like resonances [78, 79] in the transmission; whereas, in the stronger coupling case, the increased broadening of the QD energy levels smooths these resonances, resulting in more symmetrical transmission peak profiles.
@@
Figure 4: Logarithm of transmission vs. probe energy in bias range 23–24 for two strong coupling strengths: $\mathcal{C}\propto 10^{2}$ (blue) and $\mathcal{C}\propto 10^{5}$ (orange). Peaks are labeled as (a) at energy 23.127 ($\mathcal{C}\propto 10^{2}$) and (b) at energy 23.629 ($\mathcal{C}\propto 10^{5}$). Below, the corresponding normalized LDOS at these peak energies revealed how self-energy growth in strong-coupling limit disrupts the scarred patterns observed at weaker coupling, resulting in more extended wavefunction distributions.
@@
To summarize, we investigated the influence of variational scarring electron transport in two-dimensional quantum dots. In particular, we demonstrated that these scars not only survive but can also facilitate a controlled modulation of the transmission. In the weak coupling regime ($\mathcal{C}\lesssim 1$), scars enhance the transmission through the device, when oriented along the leads with a controlled perturbation, like produced by a nanotip. However, in the strong coupling limit ($\mathcal{C}\gg 1$), the effect of a individual scar is weakened by the non-scarred transport channels. Nevertheless, our study marks a step toward establishing a basis for further experimental exploration of this intriguing subclass of quantum scarring and harnessing its potential for nanoelectronic applications.
We acknowledge CSC—Finnish IT Center for Science for computational resources, and Research Council of Finland, ManyBody2D Project No. 349956, for financial support. J.K.-R. thanks the Oskar Huttunen Foundation for the financial support.

### v2

@@ Abstract
Quantum chaos – sometimes referred to as quantum chaology [1, 2] – explores the transition from quantum states to classical behavior, a topic of intense debate since the advent of quantum mechanics [3, 4]. While classical mechanics exhibits true chaos, quantum systems display signatures of chaotic behavior in distinct ways [5, 6, 7, 8], such as characteristic spectral statistics, wavefunction scarring, and other non-ergodic phenomena. At the heart of this exploration lies the correspondence principle, which is intimately linked to longstanding questions regarding measurement and decoherence in quantum mechanics [9, 10]. One proposed framework for understanding the emergence of classicality is quantum Darwinism, which posits that decoherence – by suppressing most quantum superpositions – yields environment-induced superselection (einselection). This process favors the survival of robust, pointer states that proliferate their information into the environment, while fragile quantum states are suppressed [11]. Intriguingly, these pointer states often exhibit quantum scarring, thus defying the expected ergodic behavior of quantum-chaotic systems [12].
@@
In general, open quantum dots (QDs) provide a versatile platform for investigating quantum scars, offering valuable insights into classical-quantum correspondence partially due to their mixed classical phase space structure. While a single scarred state is typically thought to be fragile against strong perturbations, such as system opening, the proliferation of scars can, however, result in significant robustness for the entire family of scarred states within a moderate energy window. Consequently, these scars can have a substantial effect on the transmission properties of quantum devices [29, 33, 43, 44]: notable transmission variations over small energy scales have been observed, with the fractal conductance fluctuations in quantum devices often attributed to their chaotic nature.
@@
In this Letter, we investigate the quantum transport of a two-dimensional (2D) device illustrated in Fig. 1 in which the coupling to the reservoirs is mediated by a pair of quantum point contacts. We assume that the QD contains impurities introducing randomly positioned bumps within the confining potential, whereas a more controlled, localized disturbance is generated via a nanotip. In particular, the perturbation gives rise to a recently discovered form of scarring within the closed QD [45, 46, 47, 48, 49, 50, 51]: these scars emerge in a synthesis of the quantum near-degeneracy in the unperturbed system and the localized nature of the perturbation. This unique birth mechanism [48] sets the variational scars apart from the conventional, Heller-type scarring [13, 17, 18], extending the concept of scarring into systems lacking suitable classical POs or when those orbits are too unstable to sustain conventional scarring. In other words, this origin allows scars to emerge and thrive under conditions that would typically inhibit conventional scarring. Furthermore, it highlights the fact that the near-degeneracies of a system can have a profound impact, not only the exact degeneracies associated with the continuous symmetries, as well as underscores the subtle role that disorder can play in a quantum device.
@@
In addition to the confining potential $V_{\textrm{ext}}$, recognizing that real quantum devices are often affected by disorder arising from impurities and imperfections – such as atoms or ions migrating into a QD system – we model the disorder potential as a sum of randomly distributed Gaussian bumps $V_{\text{imp}}=A_{n}\sum_{i}\exp(-\frac{|\bm{r}-\bm{r}_{i}|^{2}}{2\sigma_{n}^{2}})$, where $A_{n}$ and $\sigma_{n}$ represent the amplitude and width of the impurity bumps, respectively. This kind of disorder mode has been validated for a QD by density-functional theory [68, 69] and the diffusive quantum Monte-Carlo approach [62]. Moreover, we introduce a controllable localized perturbation stemming from a conducting nanotip for which a well-established approximation [70, 71, 72] is also the Gaussian profile $V_{\text{tip}}=A_{T}\exp(-{\frac{|\mathbf{r}-\mathbf{r}_{0}|^{2}}{2\sigma_{T}^{2}}})$, centered at the location $\bm{r}_{0}$, where $A_{T}$ and $\sigma_{T}$ denote its amplitude and width.
@@
As a reference, we first consider the case where there is no nanotip (i.e., $V_{tip}=0$), and the only perturbation comes from the impurity bumps $V_{\textrm{imp}}$, with amplitude $A_{n}=2$ and width $\sigma_{n}\simeq 0.05$. In this scenario, we find that several states accommodate a BB scar. Within this impurity strength regime, the perturbation $V_{\textrm{imp}}$ mostly leads to linear combinations of a single resonant set. Therefore, we can approximate the perturbed eigenstates employing the degenerate perturbation theory, i.e., by diagonalizing $V_{\textrm{imp}}$ within the near-degenerate subspace of the given resonant set. According to the variational principle, these formed states are associated with the extrema of the expectation value of the perturbation. Due to the spatially localized nature of the perturbation, the scarred states are thus favored over non-scarred states whose probability density commonly covers a much larger region of space (see, e.g., Refs. [45, 48])
@@
By the previous argument the scar orientations are mostly selected by the positions of the impurities resulting the most effective externalization of the perturbation: The BB scars shown in Fig. 2(a) orient themselves to maximize or minimize the perturbation by coinciding with or avoiding as many bumps as possible, respectively. While the scars maintain a similar alignment across a broad energy range, controlling the scar direction is challenging due to the inherent randomness of the impurity perturbation. However, by introducing an additional, stronger bump (i.e., $V_{tip}\neq 0$) via the nanotip ($A_{T}=12$ and $\sigma_{T}=0.1$), the BB scars align with the position of the nanotip, as seen in Fig. 2(b), which provides a feasible strategy to design the orientation of the scars experimentally. Furthermore, it is also worth noting that the single perturbation caused by the tip can be sufficient to generate distinct scars [46].
@@
Figure 2: (a) Examples of eigenstates in a closed, perturbed quantum dot containing random impurities (characterized by amplitude $A_{n}=2$ and width $\sigma_{n}\simeq 0.05$). The positions of impurities are represented by green dots. (b) Examples of eigenstates with the same system and quantum numbers $n$ as in (a), but now including a nanotip generating an additional, more prominent potential bump (amplitude $A_{T}=12$, width $\sigma_{T}=0.1$), which is indicated as a red dot.
Previous work [45] has demonstrated variational scars to impact the transport characteristics of a QD by enabling wavepackets to propagate through the perturbed system with higher fidelity than in the corresponding unperturbed system. This boost in fidelity stems from the fact that there can be multiple strong scars oriented similarly within a moderate energy window. [45] Moreover, these scars are robust [47, 48, 51], further underscoring their potential in a transport tool over the conventional scarring [13].
@@
Nevertheless, the influence of these scars on the electrical transmission of an open quantum device has remained elusive. To address this question, we couple the QD to two leads, effectively opening the system. The leads are modeled as semi-infinite waveguides, with harmonic confinement in the transverse direction and free propagation along the longitudinal axis. This setup ensures smooth continuity and consistent boundary conditions at the interface with the scattering region. The coupling between the leads and the QD is characterized by the rate operator $\Gamma$, which reflects the level broadening induced by the reservoirs. To quantify the coupling strength, we define a dimensionless parameter $\mathcal{C}=\textrm{Tr}[\hat{\Gamma}(\omega)]/\delta$, where $\delta$ is the mean level spacing of the isolated QD in the energy window of interest. This ratio distinguishes between weak coupling ($\mathcal{C}\ll 1$), where transmission is dominated by sharp, isolated resonances, and intermediate to strong coupling ($\mathcal{C}\gtrsim 1$), where significant level broadening leads to overlapping resonances and more delocalized transport pathways.
The transmission through the QD is then determined within the Landauer-Büttiker formalism [74, 75, 76] as
$\mathcal{T}(\omega)=\textrm{Tr}\Big{[}\hat{G}^{R}(\omega)\,\hat{\Gamma}(\omega)\,\hat{G}^{A}(\omega)\,\hat{\Gamma}(\omega)\Big{]},$ (2)
@@
$\rho(\bm{r},\omega)=-\textrm{Im}\big{[}\hat{G}^{R}(\bm{r},\omega)\big{]},$ (3)
@@
Figure 3 shows the transmission $\mathcal{T}(\omega)$ of the device as a function of energy $\omega$ in the weak coupling regime, i.e., $\mathcal{C}\ll 1$. As in the scarring analysis above, we compare two cases: one with only impurity disorder $V_{\textrm{imp}}$ and coupling strength $\mathcal{C}\simeq 0.001$, and another with a nanotip also present, $V_{\textrm{imp}}+V_{\textrm{tip}}$, and $\mathcal{C}\simeq 0.002$. Notably, we observe significantly higher transmission peaks in the nanotip scenario [peak labeled (a) in Fig. 3] compared to the noise-only instance [peak labeled (b) in Fig. 3]. We attribute this boost in transmission to the presence of BB scarring, where the noise-only case lacks the coherent pathways formed by the lead-oriented scars. This interpretation is confirmed by the LDOS, which reveals a pattern at the transmission peak energy [Fig. 3(a)] corresponding to a BB scar of the closed QD, a feature absent in the nanotip-free benchmark [Fig. 3(b)].
@@
Figure 3: Logarithm of transmission vs. probe energy within bias window 2324232423–2423 – 24 for noise-only case with $\mathcal{C}\simeq 0.001$ (blue) and nanotip+noise with $\mathcal{C}\simeq 0.002$ (orange), both in weak coupling limit ($\mathcal{C}\ll 1$). The strongest peaks are labeled as (a) for noise-only and (b) for nanotip+noise. Logarithmic scale is used due to the wide range of transmission values. Below, normalized LDOS is shown at the corresponding peak energies for (a) and (b), illustrating distinct bouncing-ball scar pattern induced by the tip. Panels (c)-(e) display normalized LDOS of nanotip+noise system at the energy corresponding to strongest transmission peak in weakly coupled leads ($\omega=23.350$), now under stronger coupling strengths $\mathcal{C}$: (c) $\simeq 7.55$, (d) $\simeq 48.34$, and (e) $\simeq 121.20$.
@@
On the other hand, panels (c)-(e) of Fig. 3 present the LDOS for the nanotip case at the same energy of 23.350, corresponding to the strongest transmission peak in the weak coupling limit. These cases are associated with three different coupling strengths, $\mathcal{C}\simeq 7.55,48.34,121.20$, respectively, from intermediate to strong coupling limit. As the coupling strength increases, the LDOS pattern that initially exhibits a clear scar becomes gradually smeared and chaotic. This reflects the fact that stronger coupling results in greater self-energy, which in turn broadens QD’s energy levels. The resulting spectral broadening diminishes the influence of individual scarred states on electron transport by delocalizing the transmission pathways.
To further explore the stronger coupling limit ($\mathcal{C}\gtrsim 1$), we focus on the same energy range where the nanotip-induced transport enhancements occur under weak coupling. Therefore, Fig. 4 shows the transmission for two cases of intermediate and strong couplings ($\mathcal{C}\simeq 12.21$ and $\mathcal{C}\simeq 96.59$). As illustrated in Fig. 4(a) and Fig. 4(b), the level broadening caused by the leads disrupts the scar patterns seen in the weak coupling regime, replacing them with more extended density distributions. Furthermore, this pronounced level broadening alters the line shape of the transmission peaks. In the weak coupling regime, the QD exhibits asymmetrical, Fano-like resonances [78, 79] in the transmission; whereas, in the stronger coupling case, the increased broadening of the QD energy levels smooths these resonances, resulting in more symmetric transmission peak profiles.
@@
Figure 4: Logarithm of transmission vs. probe energy within bias window 2324232423–2423 – 24, shown for intermediate ($\mathcal{C}\simeq 12.21$, blue), and strong ($\mathcal{C}\simeq 96.59$, orange) coupling strengths. Prominent transmissions labeled as (a) at $\omega=23.127$ and (b) at energy $\omega=23.629$. Below, the corresponding normalized LDOS at these peak energies reveals how increased self-energy disrupts the scarred patterns observed at weaker coupling, resulting in more extended wavefunction distributions.
@@
To summarize, we investigated the influence of variational scarring on electron transport in two-dimensional quantum dots. In particular, we demonstrated that these scars not only survive but can also facilitate a controlled modulation of the transmission. In the weak coupling regime ($\mathcal{C}\ll 1$), scars enhance the transmission through the device, when oriented along the leads with a controlled perturbation, such as that generated by a nanotip. However, in the intermediate to strong coupling limit ($\mathcal{C}\gtrsim 1$), the effect of an individual scar is weakened by the non-scarred transport channels. Nevertheless, our study marks a step toward establishing a basis for further experimental exploration of this intriguing subclass of quantum scarring and harnessing its potential for nanoelectronic applications.
We acknowledge CSC—Finnish IT Center for Science for computational resources, and Research Council of Finland, ManyBody2D Project No. 349956, for financial support. J.K.-R. thanks the Oskar Huttunen Foundation for the financial support. This project is also supported by the National Science Foundation (Grant No. 2403491).